OpenStudy (anonymous):

how do you convert –2x + 3y = -6 into slope intercept form? I'm getting frustrated. I have -3y = 2x +6 so far.

OpenStudy (anonymous):

Alright. You have standard form at the moment which is Ax + By + C. You need to get it into slope-intercept for which is y = mx + b. That means that you have to solve for y. -2x + 3y = -6 3y = 2x - 6 Just divided everything by 3 to get the slope-intercept form :)

OpenStudy (anonymous):

Do you need me to simplify further?

OpenStudy (anonymous):

yes please. so do i subtract 2x from both sides or something?

OpenStudy (anonymous):

Alright. THis is what I do. -2x + 3y = -6 From here. What do I do to get the -2x to the other side? Well, it's negative, so I think that I'll add 2x since -2x + 2x = 0. -2x + 2x + 3y = 2x - 6 However, remember that you have to do it to both sides to balance the equation. 3y = 2x - 6 Now, what do I do to get rid fo the 3? It's multiplied to y, so I think I'll divide it. 3y/3 = (2x - 6)/3 \(\large y = \frac{2x}{3} - 2\)

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

a little bit..

OpenStudy (anonymous):

You see, normally you would perform order of operations, PEMDAS. However, when solving for a variable, you do the opposite, SADMEP. Since there is something being added/subtracted to what you're solving for, you have to do something to get rid of that before you get to the 3y and try to get rid of the 3.

OpenStudy (anonymous):

Here's a way to make sure that it's correct. Plug in y back to the original equation. –2x + 3y = -6 –2x + 3(2x/3 - 2) = -6 -2x + 2x - 6 = -6 -6 = -6 or 0 = 0 See how both variables cancel out?

OpenStudy (anonymous):

okay i get it now :D thank you

OpenStudy (anonymous):

np :)